18-Oct-12
personnel: JohnM, TomH., KurtK/ChrisG may be coming up today to retension the main tower
Winds diminished overnight but still probably still more than 'science preferred'. Clear in am.
The Flux computer' fedora/window mgr died and flux had to be rebooted this morning.
Ingest: BlueTooth AP2 serving Stns 8-13 died last night. Gully is ingesting/archiving AP1, AP3241 for the others ok.....Investigating, see other log entry coming soon....
a17: The comms/dsm BlueTooth. May have stayed up better than before overnight. Rsync did update the data.
a2: 1m CSAT is having a diag bit.
a8: TRH's still has the 'mote \r\n' software bug. Need to fix and reprogram today.....